Solve the following second order initial value problem by the method of undetermined coefficients: y-8y +16 y = 2e, y(0)=1
Answer #1

y - syt164 = 2940 we just determine perticular soln by method of wash undetermined coefficient let, y= Axel y= 2Axent 4A Xe70- - 1 = 4 4-1 y (a) = 49000 + 4 nenor Gentzen +4 nrega 746) = 4e, tl2 3 4+4-0 6444 andyto=0) =) C2= -9 So, sole is y = et q